OpenStudy (anonymous):

For the graph of the equation y=-x^3 + 5 what is the x-intercept(s) What is the y-intercepts(s) Is the graph symmetric with respect to the x-axis? Input yes or no here : Is the graph symmetric with respect to the y-axis? Input yes or no here : Is the graph symmetric with respect to the origin? Input yes or no here : for my answers: for the x-axis (0, 3root7) for the y-axis (5,0) for the symmetry questions I have No for each of them.

OpenStudy (nottim):

x-int- Sub y=0 into the equation y-int- sub x=0 into the equation For the symmetric part, I think simply graphing it will help there. If not, remember that x^3 is to the 3rd degree

OpenStudy (nottim):

ok?

OpenStudy (nottim):

imma go

OpenStudy (anonymous):

okay. thanks

OpenStudy (anonymous):

I think my problem is that I have solved this to the best of my ability and I can't seem to get the correct answer,i'm not sure what I am doing wrong.

OpenStudy (nottim):

well, if oy uspecify, maybe I can help

OpenStudy (anonymous):

these are my answers: for my answers: for the x-axis (0, 3root7) for the y-axis (5,0) for the symmetry questions I have No for each of them.

OpenStudy (anonymous):

i let x=o and for y i got 5 and when i let y=0, it got more complicated. I ended up with 5=x^3 and i'm not sure how to either factor it out or move on to the next step.

hartnn (hartnn):

\[x^3=\sqrt5\implies x=\sqrt[3]5\]

hartnn (hartnn):

actually x^3 =5 have other 2 imaginary roots also, but you can ignore it here. just write, \((0,\sqrt[3]5)\)

hartnn (hartnn):

all your other answers are correct :)

OpenStudy (anonymous):

thank you so very much!

hartnn (hartnn):

welcome ^_^